# Audio Personas: Augmenting Social Perception via Body-Anchored Audio Cues

### v1

@@ 5.5. Measures
We measured social impression and perceived current states of others, two major uses of audio personas identified in Study #1. We collected both self-report and behavioral data. In this manuscript, we focus on discussing the self-report measures, and all behavioral measures details are provided in the supplemental materials.
To understand how audio personas shape how one forms impressions of others, we started with impression drawings and impression writings. Drawing is a common technique in psychology to measure social perception (Kamphaus and Pleiss, 1991; Cherney et al., 2006), and we used drawing to understand if participants formed impressions of others with explicit consideration of their audio persona. This was contrasted by the drawings from the control group, where the sounds existed but were not spatially associated with the person. In support of the drawings, participants also wrote down their impressions of each research assistant using at least five sentences, referred to as impression writings. Two researchers independently coded the drawings and writings to analyze if participants referenced sounds through verbal or visual representations and resolved discrepancies through discussions.
@@
To examine whether audio personas influence how people interpret the current states of others, we focused on the perceived emotions of others. Emotions were among the top-rated preferred messages participants would like to express using audio personas, as shown in Study #1. Specifically, we used the I-PANAS-SF scale (Karim et al., 2011), which includes 5 positive (active, attentive, alert, determined, inspired) and 5 negative terms (hostile, ashamed, upset, afraid, nervous), shown in Table 1.
We assessed reliability using Cronbach’s alpha for constructs with more than two items and Spearman-Brown formula for two-item constructs. Initial validation showed that reliability for likability (r) and threat perception (r) were lower than 0.7. A Principal Component Analysis identified two likability items (attractiveness, extroversion) loading in a different direction, which were dropped; this raised reliability to r. For the two-item threat perception construct, individual item results were reported, due to the inability to improve reliability via factor analysis.
@@ 6.1. Sound, Impression Management and Social Perception
In Study #1, we studied with designer-participants and identified potential use cases of audio personas to manage social impressions (e.g., personality), and signal current states (e.g., emotions). Study #2 examined the effectiveness of audio persona in influencing these dimensions from the receiver’s standpoint. Here, we discussed key insights from the findings of both studies.
Study #1 revealed that body anchoring is a key feature for audio personas, consistently present in all 48 designs created by designer-participants. Study #2 further showed that such spatial anchoring of sound shaped how social impressions were formed. Participants explicitly integrate audio personas into their impressions of others, revealed in the impression writings and drawings. This finding was contrasted with control group, where participants heard the same sounds anchored to objects and were much less likely to integrate sounds into their impressions of others. These findings align with a theoretical framework of impression formation (Livesley and Bromley, 1973; Workman and Johnson, 1991), which suggests that individuals form impressions by focusing on specific cues from the available information about a person and making inferences based on those cues. From a multisensory cue combination standpoint (Spence, 2013; Zaki, 2013), an audio persona provides both spatial and temporal alignment between an auditory cue and a person, strengthening the association of sound as available information about the person. In contrast, in the control group, sounds anchored to objects had only temporal alignment (i.e., the sound played when the person was present) but lacked spatial alignment with the person. Such a multisensory integration process could explain impression formation in the context of audio personas.
Study #1 also showed that participants leveraged different valences of sound to convey various messages. Study #2 also investigated the emotional component of audio persona. We found participants rated the person with a positive audio persona as more socially attractive, likable, and less threatening than the person with a negative audio persona. We observed a similar trend in the control group, where sound was anchored to objects during interactions with others, although the difference was not statistically significant. These results aligned with prior work that showed emotion-evoking music can influence social impressions of individuals in photographs in the same direction of the music valence (May and Hamilton, 1980; Hanser et al., 2015).
@@
While previous research suggested that sounds prime the perceived emotions of others (Logeswaran and Bhattacharya, 2009; Ignacio et al., 2021; Carroll and Young, 2005), our study did not observe a similar effect. One possible explanation is that our study put participants into more complex social settings (beyond observing static faces as in prior work), where interpreting perceived emotions of others involve processing other cues (Lange et al., 2022). Since most non-verbal cues were controlled and kept neutral in our study, participants likely perceived the primary signals as neutral, overriding the priming effect of audio. This indicates that semantic congruence (Laurienti et al., 2004; Kim and Lee, 2022) between audio personas and other contextual information might be necessary for influencing the perceived emotional states of others.
@@ 6.2. Limitations
Study #1 interviewed designer-participants to gain initial insights of audio personas usage, following prior work (Liu et al., 2023). A larger and more diverse sample, particularly participants without a technical background, is needed to understand the preferences of the broader population. Additionally, while Study #1 elicited an initial set of designs, it did not fully explore the challenges of scaling audio personas, such as in large conference settings. Future studies should invite designers to investigate techniques like audio mixing for effective deployment in large-scale, multi-user contexts.
Study #2 employed a controlled, in-lab experiment to investigate how audio personas influence the perception of others, exploring the potential uses identified in Study #1 from the receiver’s perspective. While this design provided valuable insights into how key design factors of audio personas (e.g., body-anchored, audio valence) influence social perception, it has a few limitations. First, the behavior of research assistants was pre-scripted, which ensured experimental control, but did not fully capture the nuance of everyday interactions. Future work should examine audio persona use in more socially complex settings. For instance, researchers can incorporate other communication channels (e.g., spontaneous conversations) to explore how audio personas influence impression formation with additional contextual cues. Similarly, researchers should investigate the effects of audio personas without the explicit framing of ”studying how audio impacts social interactions” to encourage naturalistic reactions. Second, the study used object-anchored sounds as the control condition, which allows us examine the exact effect of body-anchoring in audio personas. However, future work should compare audio persona with other baseline, such as no audio or non-spatialized sound. Third, in Study #2, we used only nature and animal sounds, but a wider variety of content should be explored, such as human and mechanical sounds. This study also only focused on a single style of sound characteristics, investigating variations in localization, triggers, and dynamics of audio personas will provide new insights of how audio personas influence social perception. Studying a broader range of audio persona designs will also help us to understand their differences in affecting social perception. Ultimately, the effects of audio personas in multi-user contexts and real-world social interactions beyond the lab should be explored.
@@ 6.3. Future Directions
To deploy audio personas in real-world, multi-user contexts, we lay out several future directions.
Social identities are multifaceted and adapt fluidly based on contexts (Tajfel, 1974), suggesting that users may wish to have multiple audio personas. This opens up opportunities to explore interaction techniques that enable users to seamlessly create, project, and switch between audio personas tailored to different contexts. Additionally, as users both project and receive audio personas, the control over auditory experiences becomes essential. This includes the ability to choose which audio personas to hear and to start or stop listening as desired. Future research could investigate user preferences and interface designs that offer such control, such as opt-in and opt-out mechanisms.
While audio personas were suggested to be useful in many contexts in Study #1, more investigations are needed to understand situations where audio personas should not be used. For instance, when the user needs to focus on something auditory and not be distracted by other signals, audio personas might not be well-suited. Additionally, quantitative understanding is needed to assess the effectiveness of audio personas in conveying various types of messages. This would help users leverage audio personas to efficiently communicate intended messages and avoid miscommunication that might be introduced by audio-only interactions.
As future work explores more complex audio personas, leveraging knowledge in psychoacoustics offers great potential. For example, the best human auditory localization performance occurs within - in the horizontal plane and - in elevation (Carlini et al., 2024). In Study #1, we observed that designers localized audio personas to specific body parts to create motion-specific experiences, aligning well with these perceptual thresholds. Further research could investigate computational models to support the creation of perceptually distinct audio personas. This becomes particularly important to prevent sensory confusion in multi-user contexts.
Lastly, we recognize that the value of audio personas depend on how widely they are adopted. We also acknowledge that the current headphone-based prototype poses a technology barrier, limiting the augmented social experience to users with headphones. However, audio personas do not have to be a headphone-delivered experience. For example, personal speaker-based (e.g., ”boomboxes,” (Boyer, 2014) popular in the 1980s and 90s) or clothing-integrated audio personas could achieve more universally accessible experiences. We see diverse opportunities to deploy audio personas and examine their implications for social interactions.

### v2

@@ 4.5. Measures
We adopted both self-report and behavioral measures to understand how audio personas might affect social perception.
4.5.1. Self-report measures
We started with impression drawings and impression writings. Drawing is a common technique in psychology to measure social perception (Kamphaus and Pleiss, 1991; Cherney et al., 2006), and we used drawing to understand if participants formed impressions of others with explicit consideration of their audio persona. This was contrasted by the drawings from the control group, where the sounds existed but were not spatially associated with the person. In support of the drawings, participants also wrote down their impressions of each research assistant using at least five sentences, referred to as impression writings.
@@ 4.5.1. Self-report measures
Moreover, we examined whether audio personas influence how people interpret the emotional states of others, as prior work has shown that music can shape the perceived emotions of individuals depicted in photographs. (Logeswaran and Bhattacharya, 2009; Hanser et al., 2015). Specifically, we used the I-PANAS-SF scale (Karim et al., 2011), which includes 5 positive (active, attentive, alert, determined, inspired) and 5 negative terms (hostile, ashamed, upset, afraid, nervous), shown in Table 1.
4.5.2. Behavioral measures
Two behavioral measures were collected during the icebreaker activity: interpersonal distance and seat choice. Interpersonal distance was determined by calculating the Euclidean distance between the head positions (x and z axes) of the participant and each research assistant. Head positions and rotations were continuously tracked throughout the study session using head-mounted trackers. Specifically, we focused on the minimum interpersonal distance when participants were instructed to walk toward each research assistant to present their answer sheets. This task approximates the active approach paradigms used in prior work (Tajadura-Jiménez et al., 2011). In addition, we recorded which bench participants chose to sit on at the end of the icebreaker (i.e., seat choice), with each research assistant sitting on one. The selection of a seat closer to one research assistant over the other reflects the degree of social closeness participants felt (Hayduk, 1983).
4.5.3. Overview of analysis
To analyze impression drawing and writings, two researchers independently coded the drawings and writings to analyze if participants referenced sounds through verbal or visual representations and resolved discrepancies through discussions. As exploratory measures, we analyzed the count of sound-related terms and sentiments in impression writings.
For self-report rating measures, we first assessed reliability using Cronbach’s alpha for constructs with more than two items and Spearman-Brown formula for two-item constructs. Initial validation showed that reliability for likability (r) and threat perception (r) were lower than 0.7. A Principal Component Analysis identified two likability items (attractiveness, extroversion) loading in a different direction, which were dropped; this raised reliability to r. For the two-item threat perception construct, individual item results were reported, due to the inability to improve reliability via factor analysis. We then used mixed ANOVA with audio anchoring as a between-group factor and audio valence as a within-group factor to analyze each of the self-report rating measures.
For the behavioral measures, we also used mixed ANOVA to analyze the effect on minimum interpersonal distance. For the seat choice measure, since participants chose between one of the two benches to sit, we used a chi-square test to examine the relationship between audio valence and seat choice in each condition group.
@@ 4.7. Participants
4.8. Overview of Findings
We organized the study findings into four sections. Finding #1, Inclusion of Audio in Impression Formation, examines whether participants considered audio as part of their impression of others, manifested in the impression drawings and writings. Finding #2, Impact on Social Impression, presents self-reported measures related to social impressions (social attraction, perceived threat, likability judgments), and sentiment analysis of the impression writings. Finding #3, Impact on Perceived Emotional States, includes participants’ ratings on the perceived emotional states of others. Finding #4, Impact on Social Behavior, reports participants’ seat choice selection and interpersonal distance.
@@ 4.11. Finding #3: Impact on Perceived Emotional States of Others
4.12. Finding #4: Impact on Social Behavior
Seat choice. We first analyzed the seat picking choices at the end of the icebreaker task. In the audio persona group, 14 out of 32 participants chose to sit with the research assistant that had a positive audio persona while 18 chose to sit with the research assistant with a negative audio persona. A chi-square test was conducted to examine the relationship between audio persona valence and choice of seat selection. The test statistic was , indicating that there is no statistically significant relationship between the valence of the audio persona and the participants’ choice of seating. Similarly, in the control group, no significant relationship was found between the valence of sounds anchored to the objects objects and participants’ final seat selections (). Specifically, 20 out of the 32 participants sat closer to the research assistant paired with a positive sound and 12 participants sat closer to the research assistant paired with a negative sound, with the sounds anchored to objects in the space.
Minimum interpersonal distance. Similarly, we analyzed the minimum interpersonal distance as participants approached each research assistant to show their answer sheets. This distance was calculated using the Euclidean distance between their head positions. Mixed ANOVA analysis revealed no significant main effects of either audio anchoring (F) or audio valence () on the minimum interpersonal distance. This suggests that participants generally kept a consistent distance from their partners, regardless of whether they were exposed to positive or negative sounds, and whether the sound was anchored to people or objects in the space.
4.13. Summary of Findings in Study #1
Through a controlled in-lab study, we examined the effect of audio personas on social perception. Participants exposed to audio personas (i.e., body-anchored audio cues) actively integrated these sounds into their impressions of others, as evidenced by their impression writings and drawings. In contrast, participants in the control group—who heard the same sounds anchored to objects—were much less likely to associate the sounds with the person. This contrast highlights body-anchoring as a key design feature that enables audio personas to influence social perception. Furthermore, individuals with a positive audio persona were rated as more socially attractive, likable, and less threatening than those with a negative audio persona. Lastly, we found no significant effects of audio personas on perceived emotional states of others or on behavioral measures.
5. Study #2: User Scenarios and Designs of Audio Personas
As Study #1 demonstrated the effectiveness of audio personas in influencing social perception, Study #2 explores potential use cases and design patterns for the concept. We conducted a design study with eight audio designers to uncover (1) social intentions behind using audio personas, (2) social contexts to use audio personas, and (3) design patterns of audio personas.
5.1. Methods
5.1.1. Participants
We chose to study participants with prior professional experience in audio design. This sample selection allows us to understand the potential contexts and intentions behind audio persona usage from a user perspective, as well as elicit designs with expert insights.
5.1.2. Procedure
Participants were first introduced to the concept of audio personas. They put on the prototype system, as described in Sec. 3, and interacted with an experimenter to experience three examples of audio personas: (1) a bee sound when approaching the experimenter, (2) a ”woosh” sound when the experimenter walked past, and (3) exaggerated stomping sounds from the experimenter. We chose these three demonstrations to show the technical possibilities of our prototype system, such as triggering audio personas based on proximity and anchoring sounds to body parts. These demos intended to provide inspiration, and participants were encouraged to think beyond these demo designs.
The study took approximately 75 - 90 minutes, and participants were compensated $35 for their time. The study was approved by Stanford Institutional Review Board (IRB # 71992).
5.1.3. Analysis
We video-recorded the sessions, transcribed the data, and organized it based on the guiding questions described previously. Two researchers were involved in the coding process. One researcher conducted an initial inductive thematic analysis of the transcripts to identify common themes (Braun and Clarke, 2006). The themes derived from the thematic analysis are reflected in the headings of the following subsections. Based on these themes, the two researchers independently coded each design created by the audio designers. They then discussed any discrepancies to reach consensus.
5.2. Finding #1: Social Intentions of Using Audio Personas
5.2.1. Impression management
In the designs created by participants, 21 out of 48 audio personas intended to manage impressions, focusing on influencing how others perceive users. A wide range of audio content with varying valence was used and Figure 11 shows a selection of them. For instance, P7 designed an audio persona that features auditory accessories with a jingling bell sound to reflect personal jewelry taste and personality, as shown in (Figure 11a). P5 chose to augment their footstep sounds with a trombone sound, showing the love of SpongeBob (Figure 11b). P4 designed an audio persona that emits a dog barking sound whenever they carry their dog in a pouch, signaling at a distance that they are a pet owner and cuing others to notice the dog’s presence (Figure 11e).
5.2.2. Signaling current states
5.2.3. Limitations in delivering complex information
5.2.4. Additional use cases for audio personas.
Six designs did not fall into impression management or signaling current states. Two were situated in motor learning contexts. For instance, P05 designed smoothing and relaxing sounds that played when the teacher’s and student’s body motions were aligned. In this case, the audio served as immediate feedback audible to both the sender and receiver. Two designs leveraged audio personas to evoke or recall past events. For example, P08 used the sound of ocean waves while retelling a vacation experience, immersing the listeners in the story. We also observed audio persona use in accessibility contexts. By listening to changes in one’s own audio personas during locomotion (P08), blind and low vision individuals could gain spatial awareness of the space. Last but not least, audio personas were used as motivational social gestures, allowing users to send encouragement—such as the message “you got this”—to others.
5.3. Finding #2: Social Context for Using Audio Personas
5.3.1. Intended for public and semi-public-private spaces
5.3.2. Less useful in private space
5.4. Finding #3: Design Patterns of Audio Personas
5.4.1. Localizations
We traced the sketches of audio persona locations marked by participants on the design worksheets and overlaid them to identify design patterns, as shown in Figure 14. There was variability in the specific body parts where the audio personas were attached. 25 out of 48 designs were anchored to the upper body of the users, such as to the head (n = 11), hands (n = 4), and back (n = 3). 15 designs were placed around the lower body of the users, predominantly on the feet (n = 11). In 12 of 48 designs, participants circled the whole body instead of specifying a center origin, indicating they intended the audio personas to be heard around the body without a precise origin. 8 designs placed audio personas on multiple body parts.
5.4.2. Triggers
5.4.3. Dynamics
5.5. Summary of Findings in Study #2
In Study #2, we explored social context and intended social messages for using audio personas by working with participants experienced in audio design. We identified potential use cases in managing social impressions (e.g., personality & traits, identity & belonging) and signaling current states (e.g., emotions) but are limited in communicating complex information (e.g., ideas & opinions). Participants preferred using audio personas in public (e.g., street, transport) and semi-private settings (e.g., workspace, educational institutions) to form bonds with people they do not know well, but found it less useful in private spaces. Analysis of 48 participant-generated audio personas revealed that a wide range of audio content with varied valence was used, and there are diverse design possibilities for localization, triggers, and dynamics.
@@ 6.1. Sound, Impression Management and Social Perception
Study #1 employed an in-lab, controlled study to examine the effectiveness of audio personas, body-anchored sounds, in influencing social perception.
Our findings showed that participants integrate audio personas into their impressions of others, revealed in the impression writings and drawings. In the control group, where participants heard the same sounds anchored to objects, it was much less likely to integrate sounds into their impressions of others. These findings align with a theoretical framework of impression formation (Livesley and Bromley, 1973; Workman and Johnson, 1991), which suggests that individuals form impressions by focusing on specific cues from the available information about a person and making inferences based on those cues. From a multisensory cue combination standpoint (Spence, 2013; Zaki, 2013), an audio persona provides both spatial and temporal alignment between an auditory cue and a person, strengthening the association of sound as available information about the person. In contrast, in the control group, sounds anchored to objects had only temporal alignment (i.e., the sound played when the person was present) but lacked spatial alignment with the person. Such a multisensory integration process could explain impression formation in the context of audio personas.
Study #1 also investigated the emotional component of audio persona. We found participants rated the person with a positive audio persona as more socially attractive, likable, and less threatening than the person with a negative audio persona. We observed a similar trend in the control group, where sound was anchored to objects during interactions with others, although the difference was not statistically significant. These results aligned with prior work that showed emotion-evoking music can influence social impressions of individuals in photographs in the same direction of the music valence (May and Hamilton, 1980; Hanser et al., 2015).
@@
While previous research suggested that sounds prime the perceived emotions of others (Logeswaran and Bhattacharya, 2009; Ignacio et al., 2021; Carroll and Young, 2005), our study did not observe a similar effect. One possible explanation is that our study put participants into more complex social settings (beyond observing static faces as in prior work), where interpreting perceived emotions of others involves processing other cues (Lange et al., 2022). Since most non-verbal cues were controlled and kept neutral in our study, participants likely perceived the primary signals as neutral, overriding the priming effect of audio. This indicates that semantic congruence (Laurienti et al., 2004; Kim and Lee, 2022) between audio personas and other contextual information might be necessary for influencing the perceived emotional states of others.
We did not observe main effects on certain behavioral measures, such as seat choice and interpersonal distance. These results are partially consistent with prior research on how sound influences interpersonal distance. While previous studies have found effects on audio in interpersonal distance, these were typically observed in passive approach tasks—where the experimenter approached the participant, and the participant indicated when to stop (Tajadura-Jiménez et al., 2011). In contrast, no effects were found in conditions where participants actively approached the experimenter (Tajadura-Jiménez et al., 2011). In our study, participants actively chose where to position themselves, which may explain the lack of significant behavioral effects and need further investigation.
6.2. Design Considerations for Audio Personas
Study #2 involved audio designers to elicit potential use cases and design patterns for audio personas. Based on these findings, we synthesized following design considerations.
Designing for shared spaces. Considering social contexts to use audio personas, participants in our design study preferred using audio personas in public (e.g., street, transport) and semi-private settings (e.g., workspace, educational institutions) to form bonds with people they do not know well. As audio personas are envisioned for these shared spaces, a key design challenge is how to effectively manage the projection and reception of multiple audio personas. It’s easy to imagine situations where hearing multiple audio personas simultaneously could become distracting or overwhelming. One potential solution is to draw inspiration from online social media platforms, where users manage their networks by choosing who to subscribe or unsubscribe. Similarly, audio persona systems could give users control over who can hear their audio persona and whose personas they want to receive. Additionally, as suggested by several designs in from Study #2, blending or mixing multiple personas into a single ambient layer could help users sense the overall ”vibe” of the space without being overwhelmed by different sounds. Users could then selectively tune into a specific person’s audio persona if desired.
Privacy, security and ethical considerations. Considering the social intentions behind using audio personas, participants in our design study expressed a preference for using them to manage social impressions (e.g., personality & traits, identity & belonging) and signaling current states (e.g., emotions). Building upon these use cases, security, privacy and ethical considerations need to be taken into account.
Given that audio personas may contain personal information intended for others, privacy protections are essential. Users should have full control over with whom their audio personas are shared. It is also important that users are made aware of when their audio personas are being heard during interactions—for example, through notifications. Additionally, because audio personas are tied to a user’s spatial location, continuous broadcasting may unintentionally reveal behavioral patterns that users did not intend to share. To mitigate this risk, designers might consider suggesting spatial and temporal limits—such as setting the maximum distance within which an audio persona can be heard, or maximum duration it remains active.
Designers should also be mindful of the potential harms associated with misuse of audio personas and develop appropriate intervention strategies. For instance, malicious actors could exploit the system by broadcasting loud, overwhelming sounds to capture attention or by transmitting inappropriate audio content. To mitigate such risks, one approach is to implement content moderation mechanisms, such as pre-filtering, to block disruptive or harmful audio before it reaches listeners. Additionally, in cases where audio personas are also audible to the senders themselves, potential effects on gait and movement should be considered, as highlighted by prior work on movement sonification (D’Adamo et al., 2024; Tajadura-Jiménez et al., 2015b).
Last but not least, we recognize that the value of audio personas depend on how widely they are adopted. We also acknowledge that the current headphone-based prototype poses a technology barrier, limiting the augmented social experience to users with headphones. However, audio personas do not have to be a headphone-delivered experience. For example, personal speaker-based (e.g., ”boomboxes,” (Boyer, 2014) popular in the 1980s and 90s) or clothing-integrated audio personas could achieve more universally accessible experiences.
Individual difference to audio stimuli. Analysis of 48 participant-generated designs indicates that audio personas offer rich design possibilities. Participants explored a wide range of audio content with varied emotional valence and highlighted diverse patterns for localizations, triggers, and dynamics. As designers continue to explore the design space of audio personas, one consideration to be noted is individual differences to audio stimuli. For instance, certain sounds might be deemed positive by some people while deemed as negative by others. Thus, communication through audio personas may break down if senders and receivers hold different mental models of the same sounds. A potential solution is to learn from usage patterns and construct an audio persona library that includes designs with commonly understood meanings. This is similar to emoji libraries commonly seen on social media platforms. Building on such a shared audio persona library, designers could then offer customization options to support personalized expression.
@@ 6.3. Limitations
Study #1 employed a controlled, in-lab experiment to investigate whether audio personas social perception. While this design provided valuable insights into how key design factors of audio personas (e.g., body-anchored, audio valence) influence social perception, it has a few limitations. First, the behavior of research assistants was pre-scripted, which ensured experimental control, but did not fully capture the nuance of everyday interactions. Future work should examine audio persona use in more socially complex settings. For instance, researchers can incorporate other communication channels (e.g., spontaneous conversations) to explore how audio personas influence impression formation with additional contextual cues. Similarly, researchers should investigate the effects of audio personas without the explicit framing of ”studying how audio impacts social interactions” to encourage naturalistic reactions.
Second, the study used object-anchored sounds as the control condition, which allows us to examine the exact effect of body-anchoring in audio personas. However, future work should compare audio persona with other baseline, such as no audio or non-spatialized sound. Third, in Study #1, we used only nature and animal sounds, but a wider variety of content should be explored, such as human and mechanical sounds. This study also only focused on a single style of sound characteristics, investigating variations in localization, triggers, and dynamics of audio personas will provide new insights of how audio personas influence social perception. Studying a broader range of audio persona designs will also help us to understand their differences in affecting social perception. Ultimately, the effects of audio personas in multi-user contexts and real-world social interactions beyond the lab should be explored.
Study #2 interviewed participants experienced in audio design to gain initial insights of audio personas usage, following prior work (Liu et al., 2023). A larger and more diverse sample, particularly participants without a technical background, is needed to understand the preferences of the broader population. Additionally, while Study #2 elicited an initial set of designs, it did not fully explore the challenges of scaling audio personas, such as in large conference settings. Future studies should invite designers to investigate techniques like audio mixing for effective deployment in large-scale, multi-user contexts.
@@ 6.4. Future Directions
To further explore the potential of audio personas, future research should investigate their use in more naturalistic settings. This would extend the findings from our in-lab study and build toward the potential use cases identified in the design study. One promising direction is the deployment of audio persona systems at conferences. Conferences are semi-public-private environments where attendees often seek to connect with new people over the course of a few days. Such settings offer a unique opportunity to explore the use cases and social functions of audio personas in a scalable yet manageable way. Moreover, they provide a valuable platform for studying how individuals manage multiple audio personas to present themselves differently across varying contexts (Tajfel, 1974), as well as for exploring design strategies to support hearing multiple audio personas in shared spaces.
As audio personas move beyond in-lab environments into more naturalistic settings, another direction for future work lies in exploring the psychological dynamics underlying their use. In the in-lab study, we were able to control the behavior and facial expressions of research assistants to remain neutral, isolating the effect of audio personas. However, in real-world scenarios, a person’s facial expressions and body movements are dynamic and may not always align with the tone conveyed by their audio persona. For example, someone might display a cheerful facial expression while their audio persona sounds sad. This kind of mismatch raises important research questions about the underlying processes of social perception when additional—and potentially conflicting—cues are introduced.
As future work explores more complex audio personas, leveraging knowledge in psychoacoustics offers great potential. For example, the best human auditory localization performance occurs within - in the horizontal plane and - in elevation (Carlini et al., 2024). In Study #1, we observed that audio designers localized audio personas to specific body parts to create motion-specific experiences, aligning well with these perceptual thresholds. Further research could investigate computational models to support the creation of perceptually distinct audio personas. This becomes particularly important to prevent sensory confusion in multi-user contexts.